# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

### v5

@@ 3.3 Quantum–classical instantiation details
The actor’s middle block $T$ is a 10-qubit, 10-layer PQC. Each layer applies parameterized single-qubit rotations followed by at most one CZ entangler, and all controls come from the preceding PreDNN. Concretely:
@@
All remaining architectural details222mini-batch size $N_{\text{b}}=2^{8}$,actor update frequency $N_{\text{A}}=2$., including qtDNN width ($L_{\text{h}}=2\,048$), critic topology, learning rates, and exploration noise, are listed in Table 1, where $\alpha$ is the learning rate and the Critic follows the structure from Lillicrap et al. (2019).
Table 1: hDQNN Reinforcement Learning Model Configuration for Humanoid-v4 Problem
DNN Design Hyperparameters
Components Layers Activation $N$,$M$ $d_{\text{c-link}}$ $\alpha$ $\gamma$ $\tau$
PreDNN [376,256,230] Leaky ReLU 10 $3e^{-4}$ $5e^{-3}$
qtDNN [220,2 048,10] Leaky ReLU $3e^{-4}$
PQC 10,10
PostDNN [20,256,17] Leaky ReLU 10 $3e^{-4}$ $5e^{-3}$
Critic [[376,256],[273,256,1]] Leaky ReLU $3e^{-4}$ 0.99 $5e^{-3}$
@@ 3.4 Experimental protocol
We train each agent for up to 7 20072007\,2007 200 episodes (750 000absent750000$\approx$ 750\,000≈ 750 000 timesteps) across four seeds. [^3] The horizon of $7,200$ episodes balances limited compute resources against the large number of ablation variants we evaluate. A single longer-horizon run is provided in §3.5.1.
The qtDNN uses a single hidden layer of size $L_{\text{h}}=2\,048$ and $N_{\text{qt}}=32$ updates per actor step. Other parameters are exposed in Table 1. PPO/SAC/TD3 baselines follow the best Hugging Face RL models on official Gymnasium Humanoid-v4 benchmark and consume the same number of steps to ensure a fair comparison.
Ablation.
For the ablation in §3.5.2 we keep the same total number of environment episodes and we note that there is an error rate $\sim 10^{-3}$ associated with each PQC gate. Within the same architecture, we run the experiment for an FC layer, an RBG layer, and the PQC layer across $n_{\text{qubits}}\in\{5,10\}$ and $n_{\text{shots}}\in\{100,1\,000\}$. We omit the $(n_{\text{qubits}},n_{\text{shots}})=(5\,100)$ configuration because the run got impractically slow for some seeds (see Appendix A.1 for more details on limitations). A single-qubit setting would behave almost classically, whereas circuits with more than $\approx 20$ qubits can trigger peaks above 30 GB in GPU memory, leaving too little head-room on the 40 GB GPU (cf. §3, Google Quantum AI (2025)). Likewise, $10$ shots are statistically insufficient and 10 0001000010\,00010 000 shots yield negligible variance reduction relative to their cost.
@@ 3.5.1 General performance
(a) Mean test return (10-episode rolling average).
@@
Figure 3: Learning dynamics of the hDQNN agent on Humanoid-v4. Shaded bands denote $95\%$ confidence intervals across seeds.
@@
Fig. 3 aggregates the four independent training runs across 7 20072007\,2007 200 episodes and reveals three salient trends. First, the mean test return (Fig. 3(a)) rises almost monotonically after the initial warm-up, surpassing the $10^{3}$ mark around episode 3 00030003\,0003 000 and accelerating sharply once the actor has experienced 5 50055005\,5005 500 episodes. The shaded $95\%$confidence band widens in this late phase, indicating that one or two seeds surge ahead while the slower runs continue to improve, albeit more gradually. Second, the policy loss (Fig. 3(b)) mirrors this behavior in reverse: it drops rapidly to $-100$ during the first 1 00010001\,0001 000 episodes, then decays almost linearly to $-200$ before steepening again in tandem with the return surge. The broader confidence interval after episode 5 50055005\,5005 500 signals larger gradient magnitudes and increased exploration of high-return regions. Finally, the concurrent monotonic gain in return and decline in loss across all seeds confirms that the qtDNN surrogate supplies gradients of sufficient fidelity throughout training, while the late-stage divergence illustrates the head-room the quantum-enhanced actor can exploit once it enters the high-performance regime.
Figure 4: Mean test return (10-rolling average of each actor’s best seed).
@@
Fig. 4 contrasts the best seed of each actor variant over a longer 14 0001400014\,00014 000-episode horizon. [^4] For clarity the PPO, SAC and TD3 reference values from best Hugging Face RL models (Zhao (2024, 2023a, 2023b)) on Gymnasium Humanoid-v4 are shown as horizontal dashed lines. The hDQNN-TD3 actor (solid blue) climbs steeply after episode 5 50055005\,5005 500 and saturates near a 6 011±83plus-or-minus6011836\,011$\pm$ 836 011 ± 83 return, surpassing the best seeds of classical TD3 (5 30653065\,3065 306), SAC (5 74257425\,7425 742) and PPO ($685$). Two additional aspects are noteworthy:
  • Sample-efficiency. hDQNN-TD3 reaches the 3 00030003\,0003 000-return mark about $800$ episodes earlier than its fully connected (FC) alternative, suggesting that the quantum layer helps discover high-value policies sooner.
  • Stability of convergence. The hDQNN-TD3 curve flattens into a narrow band after episode 11 0001100011\,00011 000, whereas the FC and RBG variants continue to exhibit larger oscillations, this indicates that the qtDNN surrogate supplies gradients of sufficient fidelity to stabilize late-stage policy refinement.
Overall, the figure supports our claim that, in the best case, a single well-tuned hDQNN seed can match the ceiling performance of the strongest classical baselines while requiring fewer interaction steps to get there.
@@ 3.5.2 Ablation study
Table 2 contrasts five actor variants at two training checkpoints [^5] Since the test return is noisy, we used the test return’s 10-episode rolling average across episodes to smoothen it and avoid recording misleading noise at the checkpoint.. The upper block (Episode 4 00040004\,0004 000) gives a mid-training snapshot. Already at this stage the $10$-qubit, 1 00010001\,0001 000-shot PQC attains a mean return of $678.7\pm 179.3$, and its best seed reaches $926.7$, beating the FC by $+296$ (mean-to-mean) and the RBG by $+322$.
The lower block (Episode 7 20072007\,2007 200) captures late-training behavior. Here, the $5$-qubit, 1 00010001\,0001 000-shot PQC leads with a mean return of 1 132.6±1 411.4plus-or-minus1132.61411.41\,132.6$\pm$ 1\,411.41 132.6 ± 1 411.4 and a best seed of 3 211.63211.63\,211.63 211.6, maintaining a $+512$ margin over FC and $+706$ over RBG. The larger 10-qubit variant still improves on FC by $+149$ despite its higher circuit cost.
@@
4 00040004\,0004 000 RBG 0 0 $356.5\pm 219.1$ $545.3$ $39.8$ 0
4 00040004\,0004 000 FC 0 0 $382.7\pm 167.5$ $603.6$ $27.4$ 0
4 00040004\,0004 000 PQC $010$ $470.8\pm 57.1$ $534.5$ $291.2$ 302 738302738302\,738302 738
4 00040004\,0004 000 PQC $1000$ 5 $521.3\pm 128.0$ $650.8$ $227.3$ 302 948302948302\,948302 948
4 00040004\,0004 000 PQC $0010$ $678.7\pm 179.3$ $926.7$ $307.3$ 359 952359952359\,952359 952
7 20072007\,2007 200 RBG 0 0 $426.8\pm 114.5$ $545.0$ $105.9$ 0
7 20072007\,2007 200 FC 0 0 $620.8\pm 393.8$ 1 207.81207.81\,207.81 207.8 $89.8$ 0
7 20072007\,2007 200 PQC $010$ $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694104694\,104694 104
7 20072007\,2007 200 PQC $1000$ 5 $1132.6\pm 1411.4$ 3 211.63211.63\,211.63 211.6 $737.9$ 769 867769867769\,867769 867
7 20072007\,2007 200 PQC $0010$ $768.9\pm 387.4$ 1 277.41277.41\,277.41 277.4 $700.9$ 761 498761498761\,498761 498
This consistent increase of performance between the FC and the PQC layer shows that the qtDNN is able to approximate the QL during back-propagation: the model is able to make use of the more complex representation of the QL during forward-prop and achieve a higher return, all while keeping stable gradients during back-propagation. Although not a direct experimental proof, this observation goes in the way of validating the local-tangent result of Theorem 1, i.e. $\|\nabla Q_{\theta}-\nabla Q\|\leq\varepsilon$ across the training.
@@ 3.5.3 Quantum calls
The PQC calls column from Table 2 shows how many quantum calls have been performed during the training. This number, $K$, is the same as the number of total steps the agent took during the training to explore its environment and populate the replay buffer, $\mathcal{M}$, for updating classical networks in off-policy training process. Each forward inference through the Actor Model makes one PQC call. Each PQC call to a quantum backend will return the estimated most probably bitstring from running identical PQC $S$ times (shots) and obtaining bitstrings from measuring all qubits in each shot. Therefore, hDQNN trained with qtDNN makes the minimal number ($K$, the same as the number of inference) of PQC calls and $S\times K$ PQC executions on an expensive physical QPU. Meanwhile, alternative quantum hardware compatible gradient methods like parameter shifts require $\sim I\times Q\times N_{\text{b}}\times S\times K$ additional physical QPU runs over the training process for our QL with $I=220$ dimensional input, $Q=10$ dimensional output, and a mini-batch size of $N_{\text{b}}=2^{8}$ in our cases.
@@ 4 Discussion
In this work, we proposed a novel way to efficiently train hDQNN compatible with NISQ quantum hardware, allowing for high-throughput mini-batch training using a surrogate classical neural network (qtDNN) to locally approximate and estimate gradients of the embedded quantum circuit for each mini-batch parameter update in off-policy reinforcement learning process. This method only requires expensive PQC calls in inference steps and therefore minimizes the number of expensive PQC calls in training significantly while still allows for efficient parallel mini-batch back-propagation to update classical weights. We further constructed hDQNN-TD3 model to learn complex, high-dimensional, and continuous-variable reinforcement learning tasks like Humanoid-v4 successfully with a final reward score comparable to that of models with widely used SOTA architectures except the recently proposed MEow from Chao et al. (2024). We further carried out ablation study by swapping the PQC to alternative classical layers and observed consistently better performance with hDQNN-TD3 against all comparable classical alternatives.
@@
Aknowledgements
This work has received partial funding from the European Union’s Horizon Europe research and innovation program under grant agreement No. 101084642. The work of G. K. has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
@@ A.1.1 Training hDQNN with qtDNN surrogate
The assumption that a quantum tangential deep neural network (qtDNN) can be trained as a good approximation to a given QL could only be reasonable for each mini-batch of training data (see Fig. 6) in off-policy reinforcement learning frameworks. It is because the sampled distribution of a general PQC is exponentially hard to simulate with classical computing resources Arute et al. [2019], Acharya et al. [2025].
Moreover, during the experiment, we implemented qtDNN as a 3-layer MLP and observed that training is most efficient when the hidden layer has $L_{\text{h}}=2^{N+1}$ neurons for a $N$-qubit wide PQC. As $N$-qubit spans a Hilbert space $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}\cong\mathbb{R}^{2^{N+1}}$, the experimentally observed best hidden layer’s linear space $\mathcal{L}_{\text{h}}\cong\mathbb{R}^{2^{N+1}}\cong\mathcal{H}_{N\text{q}}$ for $N=10$ in the experiment achieving a high score. Reducing the hidden linear space $\mathcal{L}_{\text{h}}$ to $\mathbb{R}^{2^{N}}$ has been observed to drastically reduce model performance in the final score. The observations indicate that physics-inspired appropriate multi-qubit state representation by the hidden layer is critical. Therefore, it is reasonable to expect, the number of neurons required for 3-layer MLP qtDNN scales according to $2^{N+1}$ with $N$ qubits in the PQC it approximates locally. We also note that the optimal 3-layer MLP design of qtDNN seems to be unexpectedly independent of PQC depth, $M$. This exponential resource scaling would limit the 3-layer MLP implementation of qtDNN to relatively narrow (in qubit count) PQC. The limitation on PQC width could further limit the ability for hDQNN + qtDNN to learn complex AI/ML tasks involving a vast number of correlated degrees of freedoms and hinder its applications in the Large Language Models (LLMs) DeepSeek-AI et al. [2025a, b] and advanced robotic controls in unpredictable real-world interactive conditions as in  NVIDIA [2025]. Future development of novel qtDNN architecture might significantly improve the scaling efficiency by tailoring the qtDNN design to respect the structure of a given PQC and bound its resource requirements with a function of $N_{\text{b}}$. For example, tensor network methods could approximate a PQC with a given structure efficiently Liu et al. [2024], Huang et al. [2021]. However, even with the potentially large resource requirements scaling exponentially with the number of qubits involved in sampling, we could argue that this limitation is only present in training and the inference data flow of a trained hDQNN is potentially exponentially more efficient as forward pass through the PQC could be carried out on physical QPUs.
@@
In terms of applying this hDQNN+qtDNN method to a variety of different RL environments, only Humanoid-v4 and the simple Pendulum-v1 were attempted due to time and resource limitations. Even though they represent two extremes of complexities in Gymnasium continous variable reinforcement learning benchmarks and achieved good results, the long PQC simulation time became a bottleneck making it challenging to extend the experimentation to other environments within the short time available to prepare the paper. However, it is straightforward to apply hDQNN + qtDNN for enhancing reinforcement learning in other useful scenarios involving both continuous and discrete observation and action spaces. Integration with the trailblazing AI models’ training and inference pipelines DeepSeek-AI et al. [2025a] could be among the topics in future researches.
In the end, hDQNN requires PQC to be executed on a physical or simulated QPU. Real-world QPUs in the near-term have noises (errors) in quantum operations they carry out. Therefore, in this work, we tested the noise tolerance of hDQNN + qtDNN method in learning Humanoid-v4. The hDQNN + qtDNN model having PQC bit-flip and phase-flip error rates of $10^{-3}$ performs on par and even potentially better than classical SOTA models in terms of rewards they achieved given a similar compute. However, a larger $10^{-2}$ error rate in physical quantum gates would lead to a final reward lower than SOTA models. On the other hand, we further noticed hDQNN + qtDNN model demonstrated faster learning with elevated error rates in quantum gates and attribute this to quantum gate noises (errors) having the same effect of larger exploration noises in classical SOTA models. We note that hyper parameters are kept the same between comparable alternative quantum and classical models presented in the main text. We further note that, as we took the most probable bitstring from sampling PQC over $S$ identical shots as the output $\mathbf{q}_{\text{o}}$ of a QL in hDQNN for a given input, $\mathbf{q}_{\text{i}}$, reducing the number of shots $S$, in each sampling for estimating $\mathbf{q}_{\text{o}}$ would increase the noises in $\mathbf{q}_{\text{o}}$. This increased sampling noise impacts the results in a similar way as errors in quantum operations.

### v6

@@ 3.3 Quantum–classical instantiation details
The actor’s middle block $T$ is a 10-qubit, 10-layer PQC. Each layer applies parameterised single-qubit rotations followed by at most one CZ entangler, and all controls come from the preceding PreDNN. Concretely:
@@
All remaining architectural details333mini-batch size $N_{\text{b}}=2^{8}$,actor update frequency $N_{\text{A}}=2$., including qtDNN width ($L_{\text{h}}=2\,048$), critic topology, learning rates, and exploration noise, are listed in App. E.1.
Noise regimes.
We consider (i) sampling noise from finite shots ($S\in\{100,1000\}$), (ii) gate noise via bit/phase‑flip channels with error rates $\sim 10^{-3}$, and (iii) exploration noise via Gaussian actions; qtDNN is trained directly on these noisy marginals; see App. F.5 for more details.
@@ 3.4 Experimental protocol
We train each agent for up to 7 2007\,200 episodes ($\approx 750\,000$ timesteps) across four seeds [^4] The horizon of $7,200$ episodes balances limited compute resources against the large number of ablation variants we evaluate. A single longer-horizon run is provided in §3.5.1..
The qtDNN uses a single hidden layer of size $L_{\text{h}}=2\,048$ and $N_{\text{qt}}=32$ updates per epoch. Other parameters are exposed in Table 2. PPO/SAC/TD3 baselines follow the best Hugging Face RL models on official Gymnasium Humanoid-v4 benchmark and consume the same number of steps to ensure a fair comparison.
Ablation protocol.
For the ablation in §3.5.2 we keep the same total number of environment episodes and we note that there is an error rate $\sim 10^{-3}$ associated with each PQC gate. To isolate the contribution of the PQC, we compare the full hDQNN agent to control variants where the middle block is replaced by a fully connected layer (FC), a random bit generator (RBG), or a zero-output layer (0L). The FC variant matches the qtDNN in width and activation, ensuring that performance differences reflect the expressive power of the quantum layer rather than architectural mismatch. We consider $n_{\text{qubits}}\in\{5,10\}$ and $n_{\text{shots}}\in\{100,1\,000\}$. We omit the $(n_{\text{qubits}},n_{\text{shots}})=(5,100)$ configuration due to unstable learning for some seeds (see Appendix A.1 for more details on limitations). A single-qubit setting would behave almost classically, whereas circuits with more than $\approx 20$ qubits can trigger peaks above 30 GB in GPU memory, leaving too little head-room on the 40 GB GPU (cf. §3, Google Quantum AI (2025)). Likewise, $10$ shots are statistically insufficient and 10 00010\,000 shots yield negligible variance reduction relative to their cost.
@@ 3.5.1 General performance
(a) Mean test return.
@@
(c) BCE loss.
Figure 3: hDQNN-TD3 on Humanoid-v4; 4 seeds; 10‑episode rolling average; 95% CI.
@@
Fig. 3 aggregates the four independent training runs across 7 2007\,200 episodes and reveals three salient trends. First, the mean test return (Fig. 3(a)) rises almost monotonically after the initial warm-up, surpassing the $10^{3}$ mark around episode 3 0003\,000 and accelerating sharply once the actor has experienced 5 5005\,500 episodes. The shaded $95\%$confidence band widens in this late phase, indicating that one or two seeds surge ahead while the slower runs continue to improve, albeit more gradually. Second, the policy loss (Fig. 3(b)) mirrors this behaviour in reverse: it drops rapidly to 100-100 during the first 1 0001\,000 episodes, then decays almost linearly to 200-200 before steepening again in tandem with the return surge. The broader confidence interval after episode 5 5005\,500 signals larger gradient magnitudes and increased exploration of high-return regions. Finally, the concurrent monotonic gain in return and decline in loss across all seeds confirms that the qtDNN surrogate supplies gradients of sufficient fidelity throughout training, while the late-stage divergence illustrates the head-room the quantum-enhanced actor can exploit once it enters the high-performance regime. Finally, Fig. 3(c) demonstrates empirical evidence of Theorem 1 by plotting the BCE loss during the training:
  • It begins with a warm-up phase: At epoch 0 the quantum-action vector $\mathbf{q}_{\text{act}}$ is almost zero, so the PQC outputs a highly imbalanced bit-string vector that the surrogate predicts with near-perfect accuracy. As exploration noise broadens the input distribution over the first ${\sim}150$ epochs, BCE rises to ${\approx}\ 0.45$.
  • Then, it reaches a stable plateau: From epoch 200 onward the loss settles in the 0.55-0.65 band and remains there for $99\%$ of the 8 000-epoch run. This is at least 13 % below the random-guess entropy of 0.693, i.e. each qubit is still predicted with ${\sim}65$ % accuracy, this taking into account that the PQC is noisy.
  • Throughout the training we notice transient spikes. Those isolated spikes occur when a mini-batch is label-skewed; they vanish within one update and do not accumulate.
This shows that the BCE loss is stable throughout the training. This suggests that the surrogate remained within the $\varepsilon$–ball required by Theorem 1 throughout training, supporting the assumption that gradient perturbations stayed within tolerance (see App. B for proof details and assumptions).
@@
Fig. 4 contrasts the best seed of each actor variant over a longer 14 00014\,000-episode horizon [^5] For clarity the PPO, SAC and TD3 reference values from best Hugging Face RL models (Zhao (2024; 2023a; 2023b)) on Gymnasium Humanoid-v4 are shown as horizontal dashed lines.. The hDQNN-TD3 actor (solid blue) climbs steeply after episode 5 5005\,500 and saturates near a $6\,011\pm 83$ return, surpassing the best seeds of classical TD3 (5 3065\,306), SAC (5 7425\,742) and PPO ($685$). Two additional aspects are noteworthy:
  • Sample-efficiency. hDQNN-TD3 reaches the 3 0003\,000-return mark about $800$ episodes earlier than its fully connected (FC) alternative, suggesting that the quantum layer helps discover high-value policies sooner.
  • Stability of convergence. The hDQNN-TD3 curve flattens into a narrow band after episode 11 00011\,000, whereas the FC and RBG variants continue to exhibit larger oscillations, this indicates that the qtDNN surrogate supplies gradients of sufficient fidelity to stabilize late-stage policy refinement.
Figure 4: Best seed return; equal step budget; horizontal lines are public PPO/SAC/TD3 references.
Overall, the figure supports our claim that, in the best case, a single well-tuned hDQNN-TD3 seed can match the ceiling performance of the strongest classical baselines while requiring fewer interaction steps to get there.
Other environments.
We tested the hDQNN-TD3 architecture on the Hopper-v4 environment (lower-dimensional, single-leg locomotion) across five seeds. Figure 5(a) shows that our model is able to reach a mean return on par with TD3’s performance on the former Hopper-v1 benchmark, namely stabilizing around a 3400 score after around 5000 epochs. Meanwhile, Figures 5(b) and 5(c) show the stability of the learning with a monotonously decreasing policy loss and a stable BCE loss throughout the training process.
(a) Mean test return.
(b) Policy loss.
(c) BCE loss.
Figure 5: hDQNN-TD3 on Hopper-v4; 4 seeds; 10‑episode rolling average; 95% CI.
@@ 3.5.2 Ablation study
Table 1 contrasts five actor variants at two training checkpoints [^6] Since the test return is noisy, we used the test return’s 10-episode rolling average across episodes to smoothen it and avoid recording misleading noise at the checkpoint.. The upper block (Episode 4 0004\,000) gives a mid-training snapshot. Already at this stage the $10$-qubit, 1 0001\,000-shot PQC attains a mean return of $678.7\pm 179.3$, and its best seed reaches $926.7$, beating the FC by +296+296 (mean-to-mean) and the RBG by +322+322.
The lower block (Episode 7 2007\,200) captures late-training behaviour. Here, the $5$-qubit, 1 0001\,000-shot PQC leads with a mean return of 1 $1.41\,132.6\pm$ 1\,411.4 and a best seed of 3 211.63\,211.6, maintaining a +512+512 margin over FC and +706+706 over RBG. The larger 10-qubit variant still improves on FC by +149+149 despite its higher circuit cost.
@@
4 0004\,000 RBG 0 0 $356.5\pm 219.1$ $545.3$ $39.8$ 0
4 0004\,000 FC 0 0 $382.7\pm 167.5$ $603.6$ $27.4$ 0
4 0004\,000 PQC $100$ 10 $470.8\pm 57.1$ $534.5$ $291.2$ 302 738302\,738
4 0004\,000 PQC $1000$ 5 $521.3\pm 128.0$ $650.8$ $227.3$ 302 948302\,948
4 0004\,000 PQC $1000$ 10 $678.7\pm 179.3$ $926.7$ $307.3$ 359 952359\,952
7 2007\,200 RBG 0 0 $426.8\pm 114.5$ $545.0$ $105.9$ 0
7 2007\,200 FC 0 0 $620.8\pm 393.8$ 1 207.81\,207.8 $89.8$ 0
7 2007\,200 PQC $100$ 10 $517.8\pm 95.0$ $659.2$ $712.2$ 694 104694\,104
7 2007\,200 PQC $1000$ 5 $\pm 1411.4 3 211.63\,211.6$ $737.9$ 769 867769\,867
7 2007\,200 PQC $1000$ 10 $768.9\pm 387.4$ 1 277.41\,277.4 $700.9$ 761 498761\,498
Across seeds, PQC outperforms the strongest FC baseline on average by +296+296 at episode $4,000$ and by +148+148 at episode $7,200$ (Table 1), confirming that the advantage is not seed-specific.
This consistent increase of performance between the FC and the PQC layer shows that the qtDNN is able to approximate the QL during back-propagation: the model is able to make use of the more complex representation of the QL during forward-prop and achieve a higher return, all while keeping stable gradients during back-propagation. Although not a direct experimental proof, this observation goes in the way of validating the local-tangent result of Theorem 1, i.e. $\|\nabla Q_{\theta}-\nabla Q\|\leq\varepsilon$ across the training. Two theoretical lenses that explain the advantage over an equally wide FC layer are summarised in App. G.
Although qtDNN is trained on each mini-batch independently, results in Table 1 show consistent performance across seeds with different replay distributions, suggesting robustness to moderate batch variability. Nonetheless, investigating further how buffer diversity affects surrogate generalisation is an important research direction, which we leave for future work.
@@ 3.5.3 Quantum calls
The PQC calls column from Table 1 shows how many quantum calls have been performed during the training. This number, $K$, is the same as the number of total steps the agent took during the training to explore its environment and populate the replay buffer, $\mathcal{M}$, for updating classical networks in off-policy training process. Each forward inference through the Actor Model makes one PQC call. Each PQC call runs the circuit for $S$ shots and returns the per‑qubit empirical marginal vector $\mathbf{q}_{\text{o}}$ computed from those $S$ bit-strings. Thus, training with qtDNN performs exactly $K$ PQC calls and $S\times K$ shot executions on an expensive physical QPU. Meanwhile, alternative quantum hardware compatible gradient methods like parameter shifts require $\sim I\times O\times N_{\text{b}}\times S\times K$ additional physical QPU runs over the training process for our QL with $I=220$ dimensional input, $O=10$ dimensional output, and a mini-batch size of $N_{\text{b}}=2^{8}$ in our cases.
Numerical efficiency.
Training with qtDNN performs exactly $K$ PQC calls and $S\!\cdot\!K$ shots (forward only), whereas parameter–shift would require $\tilde{O}(I\!\cdot\!O\!\cdot\!S\!\cdot\!N_{b}\!\cdot\!K)$ additional QPU runs. In our N=10,M=10N{=}10,M{=}10 setting (I=220I{=}220, O=10O{=}10, $N_{b}{=}2^{8}$), even a light budget with S=10S{=}10 implies $IOSN_{b}\approx 5.63\times 10^{6}$ shots per update, i.e., $\approx 3.38$s at $\sim 0.6\,\mu s$ per shot; over $7.5{\times}10^{5}$ updates ($\sim$7.2k episodes) this is $\gtrsim 2.5{\times}10^{6}$s ($\sim$29 days), while our qtDNN run finished in $\sim$700 minutes (App. A.1.2).
@@ 4 Discussion
In this work, we proposed a novel way to efficiently train hDQNN compatible with NISQ quantum hardware, allowing for high-throughput mini-batch training using a surrogate classical neural network (qtDNN) to locally approximate and estimate gradients of the embedded quantum circuit for each mini-batch parameter update in off-policy reinforcement learning process. This method only requires expensive PQC calls in inference steps and therefore minimizes the number of expensive PQC calls in training significantly while still allows for efficient parallel mini-batch back-propagation to update classical weights. We note that surrogate‑assisted training reduces quantum‑specific wall‑clock by more than an order of magnitude compared with parameter‑shift; see App. A.1.2 for a detailed cost model and lighter‑budget variants. We further constructed hDQNN-TD3 model to learn complex, high-dimensional, and continuous-variable reinforcement learning tasks like Humanoid-v4 successfully with a final reward score comparable to that of models with widely used SOTA architectures except the recently proposed MEow from Chao et al. (2024). We further carried out ablation study by swapping the PQC to alternative classical layers and observed consistently better performance with hDQNN-TD3 against all comparable classical alternatives.
@@ A.1.1 Training hDQNN with qtDNN surrogate
The assumption that a quantum tangential deep neural network (qtDNN) can be trained as a good approximation to a given QL can only be reasonable for each mini-batch of training data (see Fig. 7) in off-policy reinforcement learning frameworks. This is because the sampled distribution of a general PQC is exponentially hard to simulate with classical computing resources Arute et al. [2019], Acharya et al. [2025].
Moreover, during the experiment, we implemented qtDNN as a 3-layer MLP and observed that training is most efficient when the hidden layer spans a linear space $\mathcal{L}_{\text{h}}=\mathbb{R}^{2^{N+1}}$ with $2^{N+1}$ neurons for a $N$-qubit wide PQC. As $N$-qubit spans a Hilbert space $\mathcal{H}_{N\text{q}}\cong\mathbb{C}^{2^{N}}\cong\mathbb{R}^{2^{N+1}}\cong\mathcal{L}_{\text{h}}$, the observation indicates that physics-inspired accurate PQC state representation by the hidden layer can guide the design of optimal qtDNN. Therefore, it is reasonable to expect that the number of neurons required for 3-layer MLP qtDNN scales according to $2^{N+1}$ with $N$ qubits in the PQC it approximates locally. We also note that the optimal 3-layer MLP design of qtDNN seems to be unexpectedly independent of PQC depth, $M$. This exponential resource scaling would limit the 3-layer MLP implementation of qtDNN to relatively narrow (in qubit count) PQC. The limitation on PQC width could further limit the ability for hDQNN-TD3 to learn complex AI/ML tasks involving a vast number of correlated degrees of freedoms and hinder its applications in the Large Language Models (LLMs) DeepSeek-AI et al. [2025a, b] and advanced robotic controls in unpredictable real-world interactive conditions as in  NVIDIA [2025]. Future development of novel qtDNN architecture might significantly improve the scaling efficiency by tailoring the qtDNN design to respect the structure of a given PQC and bound its resource requirements with a function of $N_{\text{b}}$. For example, tensor network methods could approximate a PQC with a given structure efficiently Liu et al. [2024], Huang et al. [2021a]. However, even with the potentially large resource requirements scaling exponentially with the number of qubits involved in sampling, we could argue that this limitation is only present in training and the inference data flow of a trained hDQNN is potentially exponentially more efficient as forward pass through the PQC could be carried out on physical QPUs.
To counter the width scaling of the plain MLP we used in the qtDNN, techniques such as tensor-network surrogates, pruning, or symmetry-aware PQC compression could be used to significantly reduce the parameter footprint without affecting training fidelity. We view these optimisations as complementary to the main goal of demonstrating the feasibility of surrogate-assisted hybrid quantum training and leave their application as future work.
@@
In terms of applying this hDQNN-TD3 method to a variety of different RL environments, only Humanoid-v4 and the simple Hopper-v4 were attempted due to time and resource limitations. Even though they represent two extremes of complexities in Gymnasium continuous variable reinforcement learning benchmarks and achieved good results, the long PQC simulation time became a bottleneck making it challenging to extend the experimentation to other environments within the short time available to prepare the paper. However, it is straightforward to apply hDQNN-TD3 for enhancing reinforcement learning in other useful scenarios involving both continuous and discrete observation and action spaces. Integration with the trailblazing AI models’ training and inference pipelines DeepSeek-AI et al. [2025a] could be among the topics in future researches.
In the end, hDQNN requires PQC to be executed on a physical or simulated QPU. Real-world QPUs in the near-term have noises (errors) in quantum operations they carry out. Therefore, in this work, we tested the noise tolerance of hDQNN-TD3 method in learning Humanoid-v4 on top of the Gaussian action noise. The hDQNN-TD3 model having PQC bit-flip and phase-flip error rates of $10^{-3}$ perform on par and even potentially better than classical SOTA models in terms of rewards they achieved given a similar compute. However, a larger $10^{-2}$ error rate in physical quantum gates would lead to a final reward lower than SOTA models. On the other hand, we further noticed hDQNN-TD3 model demonstrated faster learning with elevated error rates in quantum gates and attribute this to quantum gate noises (errors) having the same effect of larger exploration noises in classical SOTA models. We note that hyperparameters are kept the same between comparable alternative quantum and classical models presented in the main text. We further note that, as we took the most probable bit-string from sampling PQC over $S$ identical shots as the output $\mathbf{q}_{\text{o}}$ of a QL in hDQNN for a given input, $\mathbf{q}_{\text{i}}$, reducing the number of shots $S$, in each sampling for estimating $\mathbf{q}_{\text{o}}$ would increase the noises in $\mathbf{q}_{\text{o}}$. This increased sampling noise impacts the results in a similar way as errors in quantum operations.